A motor company manufacture and sell cars and motorbikes. The cost of manufacturing x motorbikes and y cars is given by 2 2 C x y x xy y ( , ) 100 100 400 = + + . Each motorbike is sold for N$36 000-00 and each car is sold for N$180 000-00. Use Cramer’s rule to determine the number of motorbikes and the number of cars that should be manufactured and sold for a maximum profit  and determine the maximum profit max . (10 marks) 4.2 Use the Jacobian to test for functional dependence between the cost and the revenue functions in 4.1. (7 marks) 4.3 One of the stationary points of the function ( ) 4 4 2 2 f x y x y x xy y , 2 4 2 = + − + − is ( 2, 2 − ). Use the Hessian to test whether the given point is a maximum, minimum or a saddle point.


1
Expert's answer
2021-10-25T17:34:48-0400

4.1

Cost of manufacturing ,C(x,y)="100x^2+100xy+400y^2" where x is the number of motorbikes and y the number of cars.


"\therefore" Revenue,R(x,y)="36,000x+180,000y"


"\therefore" profit, "\pi" ="36,000x+180,000y-100x^2-100xy-400y^2"


For "\pi" max,"\frac{\delta\pi}{\delta x}=36,000-200x-100y=0"


"2x+y=360" .........................................(1)


"\frac{\delta\pi}{\delta y}=180,000-100x-800y=0"


"x+8y=1800" ........................................(2)


Using Gaussian elimination to eliminate x;

"(2x+16y)-(2x+y)=3600-360"


"15y=3240"


"y=216"


Putting this value of y in (2)


"x=1800-1728=72"


"\therefore" for maximum profit, "\pi" ,no of motorbikes=72 and no of cars=216.


Maximum profit =36,000(72)+180,000(216)-100(72)"^{2}" -100(72)(216)-400(216)"^{2}" ="\$20,736,000"


4.2

"\pi(x,y)=36,000x+180,000y-100x^2-100xy-400y^2"


"\frac{\delta^{2}\pi}{\delta x^{2}}=\frac{\delta}{\delta x}36,000-200x-100y=-200<0"


"\frac{\delta^{2}\pi}{\delta y^{2}}=-800"


"\frac{\delta^{2}\pi}{\delta x\delta y}=-100"


"\frac{\delta^{2}\pi}{\delta y\delta x}=-100"


Hessian, H "\begin{vmatrix}
 \frac{\delta^{2}\pi}{\delta x^{2}}& \frac{\delta^{2}\pi}{\delta y\delta x} \\
 \frac{\delta^{2}\pi}{\delta x\delta y}& \frac{\delta^{2}\pi}{\delta y^{2}}
\end{vmatrix}" ="\begin{vmatrix}
 -200 & -100 \\
 -100 & -800
\end{vmatrix}"


The above solution offers maximum profit.


4.3

To test for functional dependence between the cost and the revenue functions in 4.1 the Jacobian

= "\frac{\delta\ (C, R)}{\delta (x, y)}" should be zero.


"\frac{\delta\ (C, R)}{\delta (x, y)}=" "\begin{vmatrix}
 \frac{\delta C}{\delta x}& \frac{\delta C}{\delta y} \\
 \frac{\delta R}{\delta x}& \frac{\delta R}{\delta y}
\end{vmatrix}" ="\begin{vmatrix}
 200x+100y & 100x+800y \\
 36,000& 180,000
\end{vmatrix}"


=36000000"x" +18000000"y" -3600000"x" -28800000"y"


=32400000"x" -10800000"y"


=10800000"(3x-y)"


"\frac{\delta\ (C, R)}{\delta (x, y)}" =10800000"(3\times72-216)=0"


"\therefore" cost and revenue functions are functionality dependent.